A car moves with a speed of 45km/h. Find the distance travelled by the car in 30 minutes

Grade:7

3 Answers

Arun
6 years ago
 
Dear student
 
Speed = 45 km/hr
Means
In 1 hour or 60 kinutes it travels 45 km
Hence in 30 mins it will travel 45/2 = 22.5 Km
 
Regards
Arun (askIITians forum expert)
Ayan
5 years ago
Speed = 45 KM/HR.
1 hour = 60 mins = 45 Kms
30 mins= 45/2
= 22.5 or 22 ½ km per 30 min..............................
Satish sahu
4 years ago
Speed of car=45km/hr
Car travels distance in 60 minutes=45 km
Car travels distance in 1 minute=45/60=3/4
Car travels distance in 30 minutes=(3×30)/4
         =90/4  =22.5 km.